# Chat2Workflow: A Benchmark for Generating Executable Visual Workflows with Natural Language

### 21 Apr 2026 v1

@@ 3.3 Analysis of User Requirement Changes from Chat Interactions
Each task goes through 2 to 4 rounds of dialogue, and each round is regarded as a subtask. The workflow generation requirements for each round will be based on historical conversations, with additions, deletions or modifications made to the existing structure, which poses a further challenge to the model’s ability of complex task understanding and long-horizon instruction following.
To prevent skewed representations—as only 4 tasks contain four dialogue rounds—we illustrate the trend of the two evaluation metrics over the first three rounds for each model, where each data point is the average of three experimental runs (Figure 4). It can be intuitively observed from the table that for most models, the quality of the generated workflow shows a stable downward trend as the number of interaction rounds increases, which corroborates the challenges mentioned above. The slight improvement observed for GPT-5.2 model from the second to the third round is due to some tasks with only two interaction rounds, which affect the second round but not the third.
Furthermore, it can be observed that in most models, a common phenomenon occurs that as the rounds progress, the absolute value of the slope corresponding to the broken line gradually decreases. It might be explained as diminishing marginal returns. The workflows that can still be legally parsed and successfully executed in the second round have already undergone one screening, and they possess a certain ability to understand changes in workflow requirements. With more rounds of modifications, the rate of performance degradation would be reduced, reflecting the distinction between qualitative changes and cumulative quantitative effects.
@@ 3.4 Case Analysis
We take the StudyPlanner task of Education scenario as an example for analysis. In the first two rounds, the given instructions are as follows.
In the second round, the responses given by different models are shown in the Figure 5. Specifically, the visual workflow generated by GPT-5.2 within the Dify and Coze platforms are illustrated respectively in Figures 8 and 9 of the Appendix. 1) Kimi-K2-Instruct outputs an incorrect edge connection relationship. The node with id "5" is an "iteration" node, which internally contains a sub-workflow starting from the "iteration-start". However, there is only an inclusion relationship between them, and no edge connection relationship. This point has been emphasized in the node documentation. Therefore, the generated workflow can not really run. 2) GPT-5.2 creates a legal and executable workflow, while successfully resolving the given case. 3) There are logical inconsistencies in the content output by GLM-4.6. The node "code" and "iteration-start" are present in the workflow but not declared in <node_selection>, which implies the model’s grasp of workflow creation logic is flawed and incoherent in this instance, rendering the final workflow practically unusable.
@@ 3.5 An Error-Driven Agentic Baseline
Model Pass Rate Resolve Rate
Zero-Shot Agentic Zero-Shot Agentic
GPT-5.1 47.26 64.14 39.38 44.31
GPT-5.2 67.51 78.06 54.71 60.05
Table 2: Performance comparison of GPT-{5.1, 5.2} under Zero-Shot and the proposed Agentic settings.
To address the bottlenecks identified in Section 3.4 and beyond, we propose an error-driven agentic baseline to explore the performance upper bound of LLMs on the benchmark. Implemented via the OpenCode framework, this approach transitions from vanilla zero-shot prompting to a structured SKILL-based paradigm, explicitly incorporating task-critical guidelines, as well as rules for multi-turn interactions and variable references.
To mitigate context decay in subsequent interaction turns, we dynamically extract Variable Summaries from preceding workflow iterations as supplementary context. Crucially, we construct a robust execution loop featuring a 5-attempt retry mechanism with comprehensive structural and semantic verification. Upon failure, the framework triggers targeted Auto-Repair Modules to rectify the four most prevalent errors: code fence malformation, JSON decoding failures, topological sorting violations, and node selection inconsistencies. As shown in Table 2, the agentic framework yields substantial improvements, boosting Resolve Rates for GPT-5.1 and GPT-5.2 by absolute margins of 4.93% and 5.34%, respectively.
@@ 4.1 LLM-based Agents
The rise of LLM has greatly promoted the development of AGI, providing a capability foundation for LLM-centered AI agents. Since LLM-based agents can fully leverage the powerful task understanding capabilities of LLM, they are widely used to solve complex real-world problems with the aid of general tools Schick et al. (2023); Patil et al. (2024), collaborative frameworks Hong et al. (2024); Qian et al. (2023), and extensive API repositories Qin et al. (2024); Tang et al. (2023); Zhang et al. (2024a); Yang et al. (2024). Many prompt-based methods Wei et al. (2022); Yao et al. (2022); Zhu et al. (2026) are proven effective for improving performance. However, these agent methods or frameworks mainly focus on end-to-end effects, paying little attention to the requirements and norms of the intermediate processes in solving tasks, such as planning and reasoning paths, which is not conducive to stable and reliable execution, as well as the reproduction of results across various domains Ge et al. (2024); Xie et al. (2024); Liu et al. (2024); Yang et al. (2026); Liang et al. (2026).
@@ 4.2 Workflow and Language Agent Planning
Serving as an intermediate state, workflow provides a way to bridge the gap between task goals and the specific executable steps. On the one hand, by breaking down the task into several executable atomic steps and organizing them into execution plans based on formal logical relationships van der Aalst (1997); Dijkman et al. (2008), the workflow achieves complete transparency of the reasoning path, effectively enhancing the interpretability and reliability of the results. On the other hand, the introduction of prior organizational structures makes it possible to construct the logic of complex tasks using established patterns van der Aalst et al. (2003); Commoner et al. (1971) and robotic process automation Ivančić et al. (2019); Hofmann et al. (2020). Early workflow relied on meticulous manual design to avoid the hallucinations. However, hand-crafted workflows are time-consuming and labor-intensive. Therefore, many studies have shifted their focus to using LLM for automated workflow generation Shen et al. (2023a); Zeng et al. (2023), including iterative synthesis Zhang et al. (2024b); Li et al. (2024) and knowledge-augmented planning frameworks Ye et al. (2023); Zhu et al. (2024); Wornow et al. (2024); Huang et al. (2024). More related work can be found in Appendix C.
@@ 5 Conclusion
We introduce Chat2Workflow, a benchmark for evaluating LLMs’ ability to generate deployable visual workflows from natural language, and show that even when augmented with an agentic framework, current models remain fragile under structural constraints and requirement changes. We hope Chat2Workflow can provide both a realistic testbed and a concrete target for future research on automated workflow engineering, and will help bridge the gap between language model-driven workflow design and industrial deployment.
@@
Despite its contributions, Chat2Workflow has several limitations. First, while the high-quality dataset is manually verified , the current scale may not encompass the near-infinite variety of logic found in complex industrial business processes. Second, we simplified node interfaces to prioritize executability, which may not fully capture the intricate parameter configurations required in some real-world deployments. Finally, the current system incorporates only 20 high-frequency node types. These specific nodes were selected as they effectively cover most standard workflow scenarios; however, a wide range of valuable tools remain to be incorporated.
@@ Appendix B Experimental Settings
Appendix C More Related Work on Automated Workflow Evaluation
@@
It is crucial to evaluate the quality of the workflow created by LLM. Previous studies have tried to automate workflow evaluation in tool learning scenarios Li et al. (2023); Shen et al. (2023b) and fine-grained planning tasks Chen et al. (2024); Ye et al. (2024); Zheng et al. (2024), using metrics to assess format validity and tool consistency. But these works still have some limitations: Firstly, Most of them remain at the stage of abstract expression of the workflow Qiao et al. (2024), only checking key elements without considering actual execution effects Guo et al. (2024), structural dependencies Lal et al. (2024), or platform-specific constraints Shen et al. (2024). They may identify obvious format errors but cannot ensure usability. Secondly, these works focus on single-round evaluation, neglecting scenarios where requirements change over time.
@@
You are a strict Dify Workflow evaluator. Your task is to determine whether a workflow design is valid based only on the information provided. Do not make assumptions, infer missing nodes, or introduce any external knowledge. # Inputs I will provide four sections: 1. node_selection: A list or description of nodes selected for the workflow. It represents the nodes the designer claims will be used. 2. design_principle: Design principles or constraints for the Dify workflow. It is used to judge overall logical consistency. 3. workflow: The JSON string representation of the Dify workflow. You must parse this JSON to extract the nodes that are actually used. 4. gt_nodes: A list of mandatory (ground-truth) nodes. These nodes must be included to be considered valid. # Evaluation Rules All rules must be satisfied for the result to be true. If any rule fails, the result must be false. Rule 1: Ground-Truth Node Coverage Check whether node types in gt_nodes is a subset of node_selection. The node type only needs to appear, and there is no need to consider the frequency of occurrence. If any node type in gt_nodes is missing from node_selection, immediately return false. Do not continue further checks. Rule 2: Consistency and Exact Node Matching You must verify all of the following conditions: a. Logical Consistency: node_selection, design_principle, and workflow must be logically consistent. The workflow structure must not violate the stated design_principle. b. Exact Node Set Matching (Bidirectional Constraint): Extract the actual node set from the workflow JSON. This node set must exactly match the nodes declared in node_selection. The node type only needs to appear, and there is no need to consider the frequency of occurrence. Specifically: Nodes not declared in node_selection must not appear in workflow. Nodes declared in node_selection must appear in workflow. **Note: 1. The type of the "Template" node is "template-transform". 2. There is no need to distinguish between uppercase and lowercase letters.** # Final Decision Logic Return true only if Rule 1 and Rule 2 are fully satisfied. If any single condition fails, return false. There is no partial credit or “mostly correct” outcome. # Output Format Your response must follow this exact format and nothing else: <reason> [Explain the evaluation reasoning, stating which rules were satisfied or violated] </reason> <result> [true or false] </result>

### 26 May 2026 v2

@@ 3.3 Analysis of User Requirement Changes from Chat Interactions
Each task consists of two to four dialogue rounds. In each round, workflow requirements are iteratively modified based on historical context, presenting a significant challenge to the models’ long-horizon instruction following.
To ensure robust representations, we analyze the evaluation metrics over the first three rounds (Figure 4), omitting the fourth round due to its limited sample size. For most models, workflow quality consistently declines as interactions increase, corroborating the difficulty of handling cumulative modifications. The slight improvement observed for GPT-5.2 in the third round merely results from a changing denominator, as two-round tasks drop out of the calculation scope.
Furthermore, the degradation curve flattens over time. Rather than degrading at a uniform pace, the decline diminishes in later rounds. This indicates that workflows surviving the initial modifications act as a robust baseline. Once a model grasps the core logic in early turns, its capacity to accommodate further incremental changes stabilizes, reflecting a shift from drastic qualitative failures to minor quantitative errors.
Figure 5: Error probabilities across the four steps of the pass stage (corresponding to Algorithm 1).
Figure 6: Bad case analysis for the StudyPlanner task. We compare outputs from three representative models: Kimi-K2-Instruct produces an invalid edge connection in the iteration node; GPT-5.2 generates a correct and executable workflow; GLM-4.6 outputs logically inconsistent node declarations. It reveals common failure patterns.
@@ 3.4 Case Analysis
We take the StudyPlanner task of Education scenario as an example for analysis. The instructions for the first two rounds are in Appendix C.
Figure 6 illustrates the second-round responses from different models. Specifically, the visual workflows generated by GPT-5.2 on the Dify and Coze platforms are shown in Figures 10 and 11 of the Appendix. 1) Kimi-K2-Instruct generates an invalid edge. Node 5 is an iteration node that internally contains a sub-workflow starting with iteration-start. Despite documentation explicitly restricting their relationship to inclusion only, the model erroneously connects them with an edge, rendering the workflow unexecutable. 2) GPT-5.2 successfully resolves the case by generating a valid, executable workflow. 3) GLM-4.6 exhibits logical inconsistencies. It utilizes the code and iteration-start nodes in the final workflow without declaring them in <node_selection>, indicating a flawed and incoherent workflow creation logic.
@@ 3.5 An Error-Driven Agentic Baseline
To address the previously identified bottlenecks, we propose an error-driven agentic baseline to explore the performance upper bound of LLMs on the benchmark. Implemented via the OpenCode framework, this approach transitions from vanilla zero-shot prompting to a structured SKILL-based paradigm, explicitly incorporating task-critical guidelines, as well as rules for multi-turn interactions and variable references.
Figure 7: Performance comparison demonstrating the improvements of the proposed Agentic baseline over the Zero-Shot setting for GPT-5.1 and GPT-5.2.
To mitigate context decay in subsequent interaction turns, we dynamically extract Variable Summaries from preceding workflow iterations as supplementary context. Crucially, we construct a robust execution loop featuring a 5-attempt retry mechanism with comprehensive structural and semantic verification. Upon failure, the framework triggers targeted Auto-Repair Modules to rectify the four most prevalent errors: redundant code fences, JSON decoding failures, topological sorting violations, and node selection inconsistencies. As shown in Figure 7, the agentic framework yields substantial improvements, boosting Resolve Rate for GPT-5.1 and GPT-5.2 by absolute margins of 4.78% and 6.05%, respectively.
@@ 4.1 LLM-based Agents
LLM-based agents have demonstrated exceptional capabilities in solving complex real-world problems with the aid of general tools Schick et al. (2023); Patil et al. (2024), collaborative frameworks Hong et al. (2024); Qian et al. (2023), and extensive API repositories Qin et al. (2024); Tang et al. (2023); Zhang et al. (2024a); Yang et al. (2024). Many prompt-based methods Wei et al. (2022); Yao et al. (2022); Zhu et al. (2026) are proven effective for improving performance. However, these agent methods or frameworks mainly focus on end-to-end effects, paying little attention to the requirements and norms of the intermediate processes in solving tasks, such as planning and reasoning paths, which is not conducive to stable and reliable execution, as well as the reproduction of results across various domains Ge et al. (2024); Xie et al. (2024); Liu et al. (2024); Yang et al. (2026); Liang et al. (2026).
@@ 4.2 Workflow and Language Agent Planning
As an intermediate state, workflows bridge the gap between high-level goals and executable steps. By decomposing task into several executable atomic steps and organizing them into execution plans based on formal logical relationships van der Aalst (1997); Dijkman et al. (2008), the workflow achieves complete transparency of the reasoning path, effectively enhancing the interpretability and reliability of the results. Additionally, prior structural constraints facilitate complex task logic via established patterns van der Aalst et al. (2003); Commoner et al. (1971) and robotic process automation Ivančić et al. (2019); Hofmann et al. (2020). Although early workflows relied on labor-intensive manual design to avoid hallucinations, recent research has shifted towards automated generation using LLMs Shen et al. (2023a); Zeng et al. (2023). This includes iterative synthesis Zhang et al. (2024b); Li et al. (2024) and knowledge-augmented planning frameworks Ye et al. (2023); Zhu et al. (2024); Wornow et al. (2024); Huang et al. (2024). Appendix D provides more related work.
@@ 5 Conclusion
We introduce Chat2Workflow, a benchmark for evaluating LLMs’ ability to generate deployable visual workflows from natural language, and show that even when augmented with an agentic framework, current models remain fragile under structural constraints and requirement changes. Chat2Workflow can provide both a realistic testbed and a concrete target for future research on automated workflow engineering, and will help bridge the gap between language model-driven workflow design and industrial deployment.
@@
Despite its contributions, Chat2Workflow has several limitations. First, while the high-quality dataset is manually verified , the current scale may not encompass the near-infinite variety of logic found in complex industrial business processes. Second, we simplified node interfaces to prioritize executability, which may not fully capture the intricate parameter configurations required in some real-world deployments. Finally, the current system incorporates only 20 high-frequency node types. These specific nodes were selected as they effectively cover most standard workflow scenarios. However, a wide range of valuable tools remain to be incorporated.
@@ Appendix B Experimental Settings
Appendix C Instructions for the Studyplanner Task
The instructions for the first two rounds of the Studyplanner task are as follows:
Appendix D More Related Work on Automated Workflow Evaluation
@@
It is crucial to evaluate the quality of the workflow created by LLM. Previous studies have tried to automate workflow evaluation in tool learning scenarios Li et al. (2023); Shen et al. (2023b); Xu et al. (2025) and fine-grained planning tasks Chen et al. (2024); Ye et al. (2024); Zheng et al. (2024), using metrics to assess format validity and tool consistency. But these works still have some limitations: Firstly, Most of them remain at the stage of abstract expression of the workflow Qiao et al. (2024), only checking key elements without considering actual execution effects Guo et al. (2024), structural dependencies Lal et al. (2024), or platform-specific constraints Shen et al. (2024). They may identify obvious format errors but cannot ensure usability. Secondly, these works focus on single-round evaluation, neglecting scenarios where requirements change over time.
@@
You are a strict Dify Workflow evaluator. Your task is to determine whether a workflow design is valid based only on the information provided. Do not make assumptions, infer missing nodes, or introduce any external knowledge. # Inputs I will provide four sections: 1. node_selection: A list or description of nodes selected for the workflow. It represents the nodes the designer claims will be used. 2. design_principle: Design principles or constraints for the Dify workflow. It is used to judge overall logical consistency. 3. workflow: The JSON string representation of the Dify workflow. You must parse this JSON to extract the nodes that are actually used. 4. gt_nodes: A list of mandatory (ground-truth) nodes. These nodes must be included to be considered valid. # Evaluation Rules All rules must be satisfied for the result to be true. If any rule fails, the result must be false. Rule 1: Ground-Truth Node Coverage Check whether node types in gt_nodes is a subset of node_selection. The node type only needs to appear, and there is no need to consider the frequency of occurrence. If any node type in gt_nodes is missing from node_selection, immediately return false. Do not continue further checks. Rule 2: Consistency and Exact Node Matching You must verify all of the following conditions: a. Logical Consistency: node_selection, design_principle, and workflow must be logically consistent. The workflow structure must not violate the stated design_principle. b. Exact Node Set Matching (Bidirectional Constraint): Extract the actual node set from the workflow JSON. This node set must exactly match the nodes declared in node_selection. The node type only needs to appear, and there is no need to consider the frequency of occurrence. Specifically: Nodes not declared in node_selection must not appear in workflow. Nodes declared in node_selection must appear in workflow. **Note: 1. The type of the "Template" node is "template-transform". 2. Node matching is case-insensitive and ignores hyphens. 3. Treat aliases as identical (e.g., "$t$s" = "Text to Speech", "text2image" = "Text to Image").** # Final Decision Logic Return true only if Rule 1 and Rule 2 are fully satisfied. If any single condition fails, return false. There is no partial credit or “mostly correct” outcome. # Output Format Your response must follow this exact format and nothing else: <reason> [Explain the evaluation reasoning, stating which rules were satisfied or violated] </reason> <result> [true or false] </result>